find i0+. i'0+ and DE for t>0
0.25 7= 2 Bulesa o uc-ov
Answer #1

Que? Find hot, trot) and DE for tro, -m2H 2 301) 304) 06 0(-+) V 0-25 F - For tão for at t=0-) the circuit is in steady stateat trof - after redrawing the circuit -- 3014)= 3 because U(4)= 4 for t so Gv c = 03 606-4)=0 because Uletto for t>0 so vloge4 + R -3+ 5 = O =0 we Anow that – det cay = 025 = +fudt = 1 h Sudt iR = K = K but the value of ic ip the in equation on 0.25d